Write this equation in standard form: x squared minus 17 equals 8y plus 39
Answer by jim_thompson5910(35256)

x^2-17 = 8y + 39

x^2-17 - 39 = 8y

x^2-56 = 8y

(1/8)x^2-7 = y

y = (1/8)x^2-7